Your firm needs a machine which costs $50,000, and requires $2,000 in maintenance for each year of its five-year life. After five years, this machine will be replaced. The machine falls into the MACRS five-year class life category. Assume a tax rate of 35 percent and a discount rate of 10 percent. If this machine can be sold for $3,000 at the end of year 5, what is the after-tax salvage value?
Answer #1

As per MACRS five year class life category ,Depreciation rates are as follows :

Year Depreciation rates Accumulated depreciation rates
1 20% 20%
2 32% 20+32 = 52%
3 19.2% 52+19.2 = 71.2%
4 11.52% 71.2+11.52%= 82.72%
5 11.52% 82.72 +11.52 = 94.24%
6 5.76% 94.24+5.76 = 100%

Un-amortized Book value of machine at end of year 5 = Cost * (1-Accumulated Depreciation rate )

                                       = 50000 * (1-.9424)

                                       = 50000 * .0576

                                       = 2880

Gain on sale =Sale value -Book value

                     = 3000 -2880

                       = 120

Tax on gain =Gain on sale *tax rate

               = 120 *35%

                = 42

After tax salvage value = sale value - Tax on gain

                                       = 3000 - 42

                                      = $ 2958